Can anyone help me with ANY of these problems? Please!

1. Change 0.0555... to fraction form.
= ____

2. Simplify. The square root of 15 over the square root of 5 = _____

Solve the following problems for the roots.

3.
(the square root of 4x+5) = 3 + (the square root of 2x-6)
x= ___ and ___

4.
x^2 + 10x + 25 = 16
x= ___ and ___

5. Solve for the roots by factoring.Leave answer as an improper fraction.
(2x^2) - x = 21
x= ___ and ___

6. Solve for the roots by completing the square.
(3x^2)- 12x = 9
x= ____________

7. Solve for the roots using the quadratic formula. Leave answer as an improper fraction.
3x^2 + 11x - 4 = 0
x= ___ and ___

8. Solve for the roots using the quadratic formula.
2(6-x) = x(x+5)
x= ___________

9. Use the given equation and associated roots to perform the following exercise.
3x^2 + 11x - 4 = 0 and root1 = - 4 , root2 = (1/3)
Identify a, b, and c. (Hint: you may have to arrange the equation in the form ax2 + bx + c = 0.)
a =__ b = __ c = __

10. Use the given equation and associated roots to perform the following exercise.
3x^2 + 11x - 4 = 0 and root1 = - 4 , root2 = (1/3)
The value of the fraction below will be: -b/a
=___/____

11. Use the given equation and associated roots to perform the following exercise.
3x^2 + 11x - 4 = 0 and root1 = - 4 , root2 = (1/3)
root1 + root2 = -4 + (1/3)= ______

12. Use the given equation and associated roots to perform the following exercise.
3x^2 + 11x - 4 = 0 and root1 = - 4 , root2 = (1/3)
The sum of the roots is the same as: ___ /___

13. Use the given equation and associated roots to perform the following exercise.
3x^2 + 11x - 4 = 0 and root1 = - 4 , root2 = (1/3)
c/a = ___ /___

14. Use the given equation and associated roots to perform the following exercise.
3x^2 + 11x - 4 = 0 and root1 = - 4 , root2 = (1/3)
root1 x root2 = (-4) x (1/3) = ____

15. Use the given equation and associated roots to perform the following exercise.
3x^2 + 11x - 4 = 0 and root1 = - 4 , root2 = (1/3)
The product of the roots is the same as: ____ /____

16. Use the discriminant to determine the nature of the roots of the following equation.
y^2 - 5y - 3 = 0
Which best fits?
Double root
real and rational root
real and irrational root
imaginary root

17. Use the discriminant to determine the nature of the roots of the following equation.
x^2 + 2x + 5 = 0
Which best fits?
Double root
real and rational root
real and irrational root
imaginary root

18. Simplify the following algebraic expression, using " i" in the final answer.
The square root of 12 / the square root of -6 = the square root of ___ /___

19. Simplify the following algebraic expression.
(2 + the square root of -3 / 2) (2 - the square root of -3 / 2)
= _________

Answer by rfer(3386)

Here are the first three. Hope that gets you started.
1) 555/10000
2) square the equation=15/5=3
3) square the equation
4x+5=9+2x-6
2x=-2
x=-1